OpenStudy (czarluc):

y=(10^x-10^-x)/(10^x+10^-x) find x in terms of y

OpenStudy (anonymous):

Recall the definitions of the hyperbolic trig functions: \[\sinh x=\frac{e^x-e^{-x}}{2}~~\text{and}~~\cosh x=\frac{e^x+e^{-x}}{2}\] Notice that youhave \[\tanh x=\frac{\sinh x}{\cosh x}=\frac{e^x-e^{-x}}{e^x+e^{-x}}\] What we want to do is write \(10^x\) and \(10^{-x}\) in terms of the base \(e\), which we can do using some properties of logarithms and exponents: \[b^{f(x)}=\exp\left(\ln b^{f(x)}\right)=\exp\left(f(x)\ln b\right)=e^{\ln b\,f(x)}\] In this case, you have \(f(x)=x\) and \(f(x)=-x\) and \(b=10\). So, \[\begin{align*} y&=\frac{10^x-10^{-x}}{10^x+10^{-x}}\\\\ &=\frac{e^{x\ln 10}-e^{-x\ln 10}}{e^{x\ln 10}+e^{-x\ln 10}}\\\\ &=\tanh(x\ln 10)\\\\ &=\tanh\left(\ln10^x\right)\\\\ \tanh^{-1}y&=\ln 10^x\\\\ \exp\left(\tanh^{-1}y\right)&=10^x\\\\ \log_{10}\left[\exp\left(\tanh^{-1}y\right)\right]&=x \end{align*}\]

OpenStudy (mathmath333):

Alternate method. \(\large \begin{align} \color{black}{y=\dfrac{(10^x-10^{-x})}{(10^x+10^{-x})}\hspace{.33em}\\~\\ y+1=\dfrac{(10^x-10^{-x})}{(10^x+10^{-x})}+1\hspace{.33em}\\~\\ y+1=\dfrac{(10^x-10^{-x})+(10^x+10^{-x})}{(10^x+10^{-x})}\hspace{.33em}\\~\\ y+1=\dfrac{(2\times 10^x)}{(10^x+10^{-x})}\hspace{.33em}\\~\\ \dfrac{1}{y+1}=\dfrac{(10^x+10^{-x})}{(2\times 10^x)}\hspace{.33em}\\~\\ \dfrac{2}{y+1}=\dfrac{2(10^x+10^{-x})}{(2\times 10^x)}\hspace{.33em}\\~\\ \dfrac{2}{y+1}-1=\dfrac{2(10^x+10^{-x})}{(2\times 10^x)}-1\hspace{.33em}\\~\\ \dfrac{2-(y+1)}{y+1}=\dfrac{2(10^x+10^{-x})-(2\times 10^x)}{(2\times 10^x)}\hspace{.33em}\\~\\ \dfrac{(1-y)}{y+1}=\dfrac{2\times 10^{-x})}{(2\times 10^x)}\hspace{.33em}\\~\\ \dfrac{(1-y)}{y+1}=\dfrac{( 10^{-x})}{( 10^x)}\hspace{.33em}\\~\\ \dfrac{(1-y)}{y+1}=\dfrac{1}{( 10^{2x})}\hspace{.33em}\\~\\ 10^{2x}=\dfrac{y+1}{y-1} \hspace{.33em}\\~\\ 2x\log_{10} 10=\log_{10}\left(\dfrac{y+1}{y-1}\right) \hspace{.33em}\\~\\ 2x=\log_{10}\left(\dfrac{y+1}{y-1}\right) \hspace{.33em}\\~\\ x=\dfrac{\log_{10}\left(\dfrac{y+1}{y-1}\right)}{2} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

@mathmath333 in the 12th (I think) line, you made a slight error. The denominator in the logarithm should be \(1-y\).

OpenStudy (mathmath333):

yes thnkx for pointing ,let me correct it

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{y=\dfrac{(10^x-10^{-x})}{(10^x+10^{-x})}\hspace{.33em}\\~\\ y+1=\dfrac{(10^x-10^{-x})}{(10^x+10^{-x})}+1\hspace{.33em}\\~\\ y+1=\dfrac{(10^x-10^{-x})+(10^x+10^{-x})}{(10^x+10^{-x})}\hspace{.33em}\\~\\ y+1=\dfrac{(2\times 10^x)}{(10^x+10^{-x})}\hspace{.33em}\\~\\ \dfrac{1}{y+1}=\dfrac{(10^x+10^{-x})}{(2\times 10^x)}\hspace{.33em}\\~\\ \dfrac{2}{y+1}=\dfrac{2(10^x+10^{-x})}{(2\times 10^x)}\hspace{.33em}\\~\\ \dfrac{2}{y+1}-1=\dfrac{2(10^x+10^{-x})}{(2\times 10^x)}-1\hspace{.33em}\\~\\ \dfrac{2-(y+1)}{y+1}=\dfrac{2(10^x+10^{-x})-(2\times 10^x)}{(2\times 10^x)}\hspace{.33em}\\~\\ \dfrac{(1-y)}{y+1}=\dfrac{2\times 10^{-x})}{(2\times 10^x)}\hspace{.33em}\\~\\ \dfrac{(1-y)}{y+1}=\dfrac{( 10^{-x})}{( 10^x)}\hspace{.33em}\\~\\ \dfrac{(1-y)}{y+1}=\dfrac{1}{( 10^{2x})}\hspace{.33em}\\~\\ 10^{2x}=\dfrac{y+1}{1-y} \hspace{.33em}\\~\\ 2x\log_{10} 10=\log_{10}\left(\dfrac{y+1}{1-y}\right) \hspace{.33em}\\~\\ 2x=\log_{10}\left(\dfrac{y+1}{1-y}\right) \hspace{.33em}\\~\\ x=\dfrac{\log_{10}\left(\dfrac{y+1}{1-y}\right)}{2} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (czarluc):

Thanks for the replies!

OpenStudy (czarluc):

how to know when to add 1 or mulitply 2?